Finch Corporation produces products that it sells for $21 each. Variable costs per unit are $4, and annual fixed costs are $341,700. Finch desires to earn a profit of $79,900.

Required

  1. Use the equation method to determine the break-even point in units and dollars.

  2. Determine the sales volume in units and dollars required to earn the desired profit.
Answer #1
a.
At break-even point profit equals to 0
Sales = Variable expense + Fixed costs + Profit
(Units * Selling price ) = ( Units * Variable cost per unit ) + Fixed costs + Profit
( Units * 21 ) = ( Units * 4 ) + 341700 + 0
21Units = 4Units + 341700
21Units - 4Units = 341700
17Units = 341700
Units = 341700 / 17 20100
Break-even point in units 20100 Units
Break-even point in dollars = Break-even point in units * Selling price = 20100 * 21 422100
b.
Sales volume in units required = ( Fixed costs + Desired profit ) / ( Selling price - Variable cost per unit ) = ( 341700 + 79900 ) / ( 21 - 4 ) 24800 units
Sales volume in dollars required = Sales volume in units required * Selling price per unit = 24800 * 21 520800